# Transport properties in a model of confined granular mixtures at moderate densities

### v1

@@ I Introduction
A paradigmatic situation for studying transport properties in confined granular systems corresponds to a quasi-two-dimensional vibrating box.Olafsen and Urbach (1998); Losert, Cooper, and Gollub (1999); Prevost et al. (2004); Castillo, Mujica, and Soto (2012, 2015); Guzmán and Soto (2018); Castillo et al. (2019) In this setup, particles are confined in a box in which the vertical $z$-direction is slightly larger than the diameter of a particle. The system is driven by vertical vibrations of the box; thus, energy is injected into the particles’ vertical degrees of freedom as they collide with the top and bottom plates. This kinetic energy is subsequently dissipated and redistributed to the particles’ horizontal degrees of freedom by inter-particle collisions.
Nevertheless, describing the confined system above using a kinetic theory is quite intricate, primarily due to the restrictions imposed by confinement on the corresponding Boltzmann/Enskog collisional operator. Although some recent progress has been made by considering this approach,Mayo et al. (2022); Maynar, García Soria, and Brey (2022); Mayo et al. (2023) it is common in granular literature to start with a coarse-grained model that more effectively accounts for the effect of confinement on grain dynamics. This was the objective of the collisional model (referred to as the $\Delta$-model) proposed in Ref. Brito, Risso, and Soto, 2013. Apart from the coefficient of restitution, which characterizes the inelasticity of collisions, the $\Delta$-model introduces the factor $\Delta>0$ (which is responsible for velocity injection) to mimic the transfer of kinetic energy from the vertical degrees of freedom of granular particles to the horizontal ones.
@@
Since the expressions of the diffusion transport coefficients are available, as an application we have also obtained the so-called thermal diffusion factor $\Lambda$.Kincaid, Cohen, and López de Haro (1987) Knowledge of this quantity provides a segregation criterion showing the transition between regions of the parameter space where $\Lambda>0$ (where larger particles tend to accumulate near the cooler plate) to regions where $\Lambda<0$ (where larger particles tend to accumulate near the hotter plate). Since the origin of segregation is due to both gravity and the thermal gradient, we will analyze limiting cases in which the influence of each on segregation can be disentangled.
@@ II.2 Enskog kinetic equation for granular mixtures
$J_{ij}[\mathbf{r},\mathbf{v}_{1}|f_{i},f_{j}]=\sigma_{ij}^{d-1}\int\mathrm{d}{\bf v}_{2}\int\mathrm{d}\widehat{\boldsymbol{\sigma}}\;\Theta(-\widehat{{\boldsymbol{\sigma}}}\cdot{\bf g}_{12}-2\Delta_{ij})(-\widehat{\boldsymbol{\sigma}}\cdot{\bf g}_{12}-2\Delta_{ij})\alpha_{ij}^{-2}\chi_{ij}(\mathbf{r},\mathbf{r}+{\boldsymbol{\sigma}}_{ij})f_{i}(\mathbf{r},\mathbf{v}_{1}^{\prime\prime};t)$
$\times f_{j}(\mathbf{r}+{\boldsymbol{\sigma}}_{ij},\mathbf{v}_{2}^{\prime\prime};t)-\sigma_{ij}^{d-1}\int\mathrm{d}{\bf v}_{2}\int\mathrm{d}\widehat{\boldsymbol{\sigma}}\;\Theta(\widehat{{\boldsymbol{\sigma}}}\cdot{\bf g}_{12})(\widehat{\boldsymbol{\sigma}}\cdot{\bf g}_{12})\chi_{ij}(\mathbf{r},\mathbf{r}+{\boldsymbol{\sigma}}_{ij})f_{i}(\mathbf{r},\mathbf{v}_{1};t)f_{j}(\mathbf{r}+{\boldsymbol{\sigma}}_{ij},\mathbf{v}_{2};t).$ (9)
@@
$D_{t}n_{i}+n_{i}\nabla\cdot{\bf U}+\frac{\nabla\cdot{\bf j}_{i}}{m_{i}}=0,$ (11)
$D_{t}{\bf U}+\rho^{-1}\nabla\cdot\mathsf{P}=\mathbf{g},$ (12)
$D_{t}T-\frac{T}{n}\sum_{i=1}^{s}\frac{\nabla\cdot{\bf j}_{i}}{m_{i}}+\frac{2}{dn}\left(\nabla\cdot{\bf q}+\mathsf{P}:\nabla{\bf U}\right)=-\zeta\,T.$ (13)
@@
is the (global) granular temperature. In addition, $D_{t}=\partial_{t}+\mathbf{U}\cdot\nabla$ is the material derivative, $\rho=\sum_{i}\rho_{i}=\sum_{i}m_{i}n_{i}$ is the total mass density, and ${\bf V}={\bf v}-{\bf U}$ is the peculiar velocity. Apart from the granular temperature $T$, at a kinetic level it is convenient to introduce the partial temperatures $T_{i}$ for each species; they measure their mean kinetic energies. They are defined as
@@
$\times f_{ij}\left[\mathbf{r}-\lambda{\boldsymbol{\sigma}}_{ij},\mathbf{r}+(1-\lambda){\boldsymbol{\sigma}}_{ij},\mathbf{v}_{1},\mathbf{v}_{2},t\right]$
@@
$\times f_{ij}\left[\mathbf{r}-\lambda{\boldsymbol{\sigma}}_{ij},\mathbf{r}+(1-\lambda){\boldsymbol{\sigma}}_{ij},\mathbf{v}_{1},\mathbf{v}_{2},t\right].$
@@ III Chapman–Enskog method
One also expects the existence of the above two stages for granular gases. However, in the kinetic stage, the distribution function generally relaxes toward a time-dependent, nonequilibrium distribution (the homogeneous cooling state in the conventional IHS model) rather than a local equilibrium distribution. A crucial point is that, although the granular temperature $T$ is not conserved (since collisions between particles are inelastic), $T$ can still be considered a slow field, as in conventional fluids. This assumption is clearly supported by the good agreement found between granular hydrodynamics and computer simulations in several non-equilibrium situations.Dahl et al. (2002); Garzó (2019) More details on applying the Chapman–Enskog method to granular mixtures can be found in Ref. Garzó, 2019, for example.
@@ III.1 Zeroth-order approximation
$J_{ij}^{(0)}[\mathbf{v}_{1}|f_{i}^{(0)},f_{j}^{(0)}]\equiv\sigma_{ij}^{d-1}\chi_{ij}\int\mathrm{d}{\bf v}_{2}\int\mathrm{d}\widehat{\boldsymbol{\sigma}}\;$ (36)
$\times\Theta(-\widehat{{\boldsymbol{\sigma}}}\cdot{\bf g}_{12}-2\Delta_{ij})(-\widehat{\boldsymbol{\sigma}}\cdot{\bf g}_{12}-2\Delta_{ij})\alpha_{ij}^{-2}f_{i}^{(0)}(\mathbf{v}_{1}^{\prime\prime})$
$\times f_{j}^{(0)}(\mathbf{v}_{2}^{\prime\prime})-\sigma_{ij}^{d-1}\chi_{ij}\int\mathrm{d}{\bf v}_{2}\int\mathrm{d}\widehat{\boldsymbol{\sigma}}\;\Theta(\widehat{{\boldsymbol{\sigma}}}\cdot{\bf g}_{12})$
$(\widehat{\boldsymbol{\sigma}}\cdot{\bf g}_{12})f_{i}^{(0)}(\mathbf{v}_{1})f_{j}^{(0)}(\mathbf{v}_{2}).$
@@ III.1.1 Steady solution: Maxwellian approximation
A good estimate of the zeroth-order contributions to the cooling rates and the hydrostatic pressure can be obtained by assuming Maxwellian distributions $f_{i,\text{M}}(\mathbf{V})$ for the distributions $f_{i}^{(0)}(\mathbf{V})$:
$f_{i,\text{M}}(\mathbf{V})=n_{i}\left(\frac{m_{i}}{2\pi T_{i}^{(0)}}\right)^{d/2}\exp\left(-\frac{m_{i}V^{2}}{2T_{i}^{(0)}}\right).$ (49)
In this approximation, the (dimensionless) cooling rate $\zeta_{0}^{*}=\zeta^{(0)}/\nu=\sum_{i}x_{i}\gamma_{i}\zeta_{i}^{*}$ where the (dimensionless) partial cooling rates $\zeta_{i}^{*}=\zeta_{i}^{(0)}/\nu$ are given by Brito, Soto, and Garzó (2020)
@@
$-\frac{4\pi^{d/2}}{d\Gamma\left(\frac{d}{2}\right)}\sum_{j=1}^{s}x_{j}\chi_{ij}\left(\frac{\sigma_{ij}}{\overline{\sigma}}\right)^{d-1}\mu_{ji}\Delta_{ij}^{*}\left[\frac{2\mu_{ji}\Delta_{ij}^{*}}{\sqrt{\pi}}\theta_{i}^{1/2}\left(1+\theta_{ij}\right)^{1/2}-1+\mu_{ji}(1+\alpha_{ij})\left(1+\theta_{ij}\right)\right].$
Here, $\theta_{ij}=m_{i}T_{j}^{(0)}/m_{j}T_{i}^{(0)}$ gives the ratio between the mean-square velocity of the particles of the species $j$ relative to that of the particles of the species $i$. The (reduced) pressure $p^{*}$ can be finally written as
@@
$+\frac{2}{\sqrt{\pi}}\frac{m_{i}}{\overline{m}}\Delta_{ij}^{*}\left(\frac{\theta_{i}+\theta_{j}}{\theta_{i}\theta_{j}}\right)^{1/2}\Bigg],$
where $\theta_{i}=m_{i}/(\overline{m}\gamma_{i})$.
@@ IV.2 Pressure tensor
Here,
$\varphi_{i,\text{M}}(\mathbf{c})=\pi^{-d/2}\theta_{i}^{d/2}e^{-\theta_{i}c^{2}}$ (74)
is the Maxwellian approximation to the scaled distribution $\varphi_{i}(\mathbf{c})$ of the species $i$.
@@ V.3 Mechanically equivalent particles
$\int\mathrm{d}\mathbf{V}\;R_{i,\lambda\beta}(\mathbf{V})\mathcal{K}_{ij,\lambda}\left[\frac{\partial f_{j}^{(0)}}{\partial V_{\beta}}\right]=2^{d}(d-1)x_{i}x_{j}\chi\phi\Bigg\{\frac{1}{4}(1+\alpha)(1-3\alpha)-\Delta^{*}\Bigg[\sqrt{\frac{2}{\pi}}(1+2\alpha)+\Delta^{*}\Bigg]\Bigg\},$ (103)
@@
where [^1] A misprint in the calculation of the collision integral (97) for monocomponent granular gases was found in Ref. Garzó, Brito, and Soto, 2018 while the present paper was written. The expression (103) corrects this result. For this reason, the expression (106) for $\eta_{k}^{*}$ differs slightly from the one derived in Ref. Garzó, Brito, and Soto, 2018.

### v2

@@ I Introduction
A paradigmatic situation for studying transport properties in confined granular systems corresponds to a quasi-two-dimensional vibrating box.Olafsen and Urbach (1998); Losert, Cooper, and Gollub (1999); Prevost et al. (2004); Castillo, Mujica, and Soto (2012, 2015); Guzmán and Soto (2018); Castillo et al. (2019) In this setup, particles are confined in a box in which the vertical $z$-direction is slightly larger than the diameter of a particle. The system is driven by vertical vibrations of the box; thus, energy is injected into the particles’ vertical degrees of freedom as they collide with the bottom plate of the box. This kinetic energy is subsequently dissipated and redistributed to the particles’ horizontal degrees of freedom by inter-particle collisions.
Figure 1: Schematic illustration of the confined quasi-two-dimensional system. Vertical vibration is imposed on the box to inject external energy into the system. During nonplanar collisions between grains, the vertical energy gained from the vibration of the plate is transferred to the $xy$ components of the velocities. This kinetic energy is then dissipated and redistributed among the particles’ horizontal degrees of freedom due to their collisions. It is important to note that, although our study aims to capture the phenomenology of confined systems, we are modeling an unconfined two-dimensional system where collisions are described by the $\Delta$-model.
Nevertheless, describing the confined system above using a kinetic theory is quite intricate, primarily due to the restrictions imposed by confinement on the corresponding Boltzmann/Enskog collisional operator. Although some recent progress has been made by considering this approach,Mayo et al. (2022); Maynar, García Soria, and Brey (2022); Mayo et al. (2023) it is common in granular literature to start with a coarse-grained model that more effectively accounts for the effect of confinement on grain dynamics. This was the objective of the collisional model (referred to as the $\Delta$-model) proposed in Ref. Brito, Risso, and Soto, 2013. Apart from the coefficient of restitution, which characterizes the inelasticity of collisions, the $\Delta$-model introduces the factor $\Delta>0$ (which is responsible for velocity injection) to mimic the transfer of kinetic energy from the vertical degrees of freedom of granular particles to the horizontal ones. A schematic illustration of the studied system is provided in Fig. 1.
@@
Since the expressions of the diffusion transport coefficients are available, as an application we have also obtained the so-called thermal diffusion factor $\Lambda$.Kincaid, Cohen, and López de Haro (1987) Knowledge of this quantity provides a segregation criterion showing the transition between regions of the parameter space where $\Lambda>0$ (where larger particles tend to accumulate near the cooler plate) to regions where $\Lambda<0$ (where larger particles tend to accumulate near the hotter plate). The first situation ($\Lambda>0$) is usually referred to in granular literature as the Brazil nut Effect (BNE), while the second situation ($\Lambda<0$) is referred to as the Reverse Brazil nut Effect (BNE). Since the origin of segregation is due to both gravity and the thermal gradient, we will analyze limiting cases in which the influence of each on segregation can be disentangled.
@@ II.2 Enskog kinetic equation for granular mixtures
$J_{ij}[\mathbf{r},\mathbf{v}_{1}|f_{i},f_{j}]=\sigma_{ij}^{d-1}\int\mathrm{d}{\bf v}_{2}\int\mathrm{d}\widehat{\boldsymbol{\sigma}}\;\Theta(-\widehat{{\boldsymbol{\sigma}}}\cdot{\bf g}_{12}-2\Delta_{ij})(-\widehat{\boldsymbol{\sigma}}\cdot{\bf g}_{12}-2\Delta_{ij})\alpha_{ij}^{-2}\chi_{ij}(\mathbf{r},\mathbf{r}+{\boldsymbol{\sigma}}_{ij})f_{i}(\mathbf{r},\mathbf{v}_{1}^{\prime\prime};t)$ (9)
$\times f_{j}(\mathbf{r}+{\boldsymbol{\sigma}}_{ij},\mathbf{v}_{2}^{\prime\prime};t)-\sigma_{ij}^{d-1}\int\mathrm{d}{\bf v}_{2}\int\mathrm{d}\widehat{\boldsymbol{\sigma}}\;\Theta(\widehat{{\boldsymbol{\sigma}}}\cdot{\bf g}_{12})(\widehat{\boldsymbol{\sigma}}\cdot{\bf g}_{12})\chi_{ij}(\mathbf{r},\mathbf{r}+{\boldsymbol{\sigma}}_{ij})f_{i}(\mathbf{r},\mathbf{v}_{1};t)f_{j}(\mathbf{r}+{\boldsymbol{\sigma}}_{ij},\mathbf{v}_{2};t).$
@@
$\text{D}_{t}n_{i}+n_{i}\nabla\cdot{\bf U}+\frac{\nabla\cdot{\bf j}_{i}}{m_{i}}=0,$ (11)
$\text{D}_{t}{\bf U}+\rho^{-1}\nabla\cdot\mathsf{P}=\mathbf{g},$ (12)
$\text{D}_{t}T-\frac{T}{n}\sum_{i=1}^{s}\frac{\nabla\cdot{\bf j}_{i}}{m_{i}}+\frac{2}{dn}\left(\nabla\cdot{\bf q}+\mathsf{P}:\nabla{\bf U}\right)=-\zeta\,T.$ (13)
@@
is the (global) granular temperature. In addition, $\text{D}_{t}=\partial_{t}+\mathbf{U}\cdot\nabla$ is the material derivative, $\rho=\sum_{i}\rho_{i}=\sum_{i}m_{i}n_{i}$ is the total mass density, and ${\bf V}={\bf v}-{\bf U}$ is the peculiar velocity. Apart from the granular temperature $T$, at a kinetic level it is convenient to introduce the partial temperatures $T_{i}$ for each species; they measure their mean kinetic energies. They are defined as
@@
$\times f_{ij}\left[\mathbf{r}-\lambda{\boldsymbol{\sigma}}_{ij},\mathbf{v}_{1},\mathbf{r}+(1-\lambda){\boldsymbol{\sigma}}_{ij},\mathbf{v}_{2},t\right]$
@@
$\times f_{ij}\left[\mathbf{r}-\lambda{\boldsymbol{\sigma}}_{ij},\mathbf{v}_{1},\mathbf{r}+(1-\lambda){\boldsymbol{\sigma}}_{ij},\mathbf{v}_{2},t\right].$
@@ III Chapman–Enskog method
One also expects the existence of the above two stages for granular gases. However, in the kinetic stage, the distribution function generally relaxes toward a time-dependent, nonequilibrium distribution (the homogeneous cooling state in the conventional IHS model) rather than a local equilibrium distribution. A crucial point is that, although the kinetic energy is not conserved (since collisions between particles are inelastic), the granular temperature $T$ can still be considered a slow field, as in conventional fluids. This assumption is clearly supported by the good agreement found between granular hydrodynamics and computer simulations in several non-equilibrium situations.Dahl et al. (2002); Garzó (2019) More details on applying the Chapman–Enskog method to granular mixtures can be found in Ref. Garzó, 2019, for example.
@@
What follows uses Latin indices to label the particle species (running from 1 to $s$) and Greek indices to label the spatial dimensions ($d=2$ for disks and $d=3$ for spheres). In addition, Einstein summation convention over repeated Greek indices is assumed in this paper.
@@ III.1 Zeroth-order approximation
$J_{ij}^{(0)}[\mathbf{v}_{1}|f_{i}^{(0)},f_{j}^{(0)}]$ $\equiv$ $\sigma_{ij}^{d-1}\chi_{ij}\int\mathrm{d}{\bf v}_{2}\int\mathrm{d}\widehat{\boldsymbol{\sigma}}\;\Theta(-\widehat{{\boldsymbol{\sigma}}}\cdot{\bf g}_{12}-2\Delta_{ij})(-\widehat{\boldsymbol{\sigma}}\cdot{\bf g}_{12}-2\Delta_{ij})\alpha_{ij}^{-2}f_{i}^{(0)}(\mathbf{v}_{1}^{\prime\prime})f_{j}^{(0)}(\mathbf{v}_{2}^{\prime\prime})$ (36)
$-\sigma_{ij}^{d-1}\chi_{ij}\int\mathrm{d}{\bf v}_{2}\int\mathrm{d}\widehat{\boldsymbol{\sigma}}\;\Theta(\widehat{{\boldsymbol{\sigma}}}\cdot{\bf g}_{12})(\widehat{\boldsymbol{\sigma}}\cdot{\bf g}_{12})f_{i}^{(0)}(\mathbf{v}_{1})f_{j}^{(0)}(\mathbf{v}_{2}).$
@@ III.1.1 Steady solution: Maxwellian approximation
According to Eq. (41), determining the zeroth-order contributions to the cooling rates requires knowledge of the distribution functions $f_{i}^{(0)}(\mathbf{V})$. In the steady state and for elastic collisions, $\gamma_{i}=1$ and $\Delta_{ij}=0$ and the solution to Eq. (46) is the Gaussian or Maxwellian distribution $\varphi_{1}(\mathbf{c})=\varphi_{2}(\mathbf{c})=\cdots=\varphi_{s}(\mathbf{c})=\pi^{-d/2}e^{-c^{2}}$. However, an exact solution to Eq. (46) has not yet been obtained for inelastic collisions. A systematic approximation for the isotropic distributions $\varphi_{i}(\mathbf{c})$ can be found by expanding them into a complete set of orthogonal polynomials with a Gaussian measure. In practice, generalized Laguerre or Sonine polynomials $S_{p}^{(i)}(c^{2})$ are used in kinetic theory.Chapman and Cowling (1970); Ferziger and Kaper (1972) Thus, $\varphi_{i}(\mathbf{c})$ can be written as
$\varphi_{i}(\mathbf{c})=\varphi_{i,\text{M}}(\mathbf{c})\Big[1+\sum_{p=1}^{\infty}a_{p}^{(i)}S_{p}^{(i)}(c^{2})\Big],$ (49)
where the cumulants $a_{p}^{(i)}$ are given in terms of the velocity moments of $\varphi_{i}$. However, results obtained for binary granular mixturesBrito, Soto, and Garzó (2020) have clearly shown that in general the magnitude of the coefficients $a_{p}^{(i)}$ is very small. For practical purposes, one can therefore take the Maxwellian approximations
$\varphi_{i,\text{M}}(\mathbf{c})=\pi^{-d/2}\theta_{i}^{d/2}e^{-\theta_{i}c^{2}},$ (50)
or equivalently
$f_{i,\text{M}}(\mathbf{V})=n_{i}\left(\frac{m_{i}}{2\pi T_{i}^{(0)}}\right)^{d/2}\exp\left(-\frac{m_{i}V^{2}}{2T_{i}^{(0)}}\right)$ (51)
to estimate the zeroth-order contributions to the cooling rates. Here, $\theta_{i}=m_{i}/(\overline{m}\gamma_{i})$. In this approximation, the (dimensionless) cooling rate $\zeta_{0}^{*}=\zeta^{(0)}/\nu=\sum_{i}x_{i}\gamma_{i}\zeta_{i}^{*}$ where the (dimensionless) partial cooling rates $\zeta_{i}^{*}=\zeta_{i}^{(0)}/\nu$ are given by Brito, Soto, and Garzó (2020)
@@
$-\frac{4\pi^{d/2}}{d\Gamma\left(\frac{d}{2}\right)}\sum_{j=1}^{s}x_{j}\chi_{ij}\left(\frac{\sigma_{ij}}{\overline{\sigma}}\right)^{d-1}\mu_{ji}\Delta_{ij}^{*}\left[\frac{2\mu_{ji}\Delta_{ij}^{*}}{\sqrt{\pi}}\theta_{i}^{1/2}\left(1+\theta_{ij}\right)^{1/2}-1+\mu_{ji}(1+\alpha_{ij})\left(1+\theta_{ij}\right)\right],$
where $\theta_{ij}=m_{i}T_{j}^{(0)}/m_{j}T_{i}^{(0)}$ gives the ratio between the mean-square velocity of the particles of the species $j$ relative to that of the particles of the species $i$. In addition, the (reduced) pressure $p^{*}$ can be finally written as
@@
$+\frac{2}{\sqrt{\pi}}\frac{m_{i}}{\overline{m}}\Delta_{ij}^{*}\left(\frac{\theta_{i}+\theta_{j}}{\theta_{i}\theta_{j}}\right)^{1/2}\Bigg].$
@@ V.3 Mechanically equivalent particles
$\int\mathrm{d}\mathbf{V}\;R_{i,\lambda\beta}(\mathbf{V})\mathcal{K}_{ij,\lambda}\left[\frac{\partial f_{j}^{(0)}}{\partial V_{\beta}}\right]=2^{d}(d-1)x_{i}x_{j}\chi\phi nT\Bigg\{\frac{1}{4}(1+\alpha)(1-3\alpha)-\Delta^{*}\Bigg[\sqrt{\frac{2}{\pi}}(1+2\alpha)+\Delta^{*}\Bigg]\Bigg\},$ (104)
@@
whereNote (1)
@@
The (reduced) cooling rate $\zeta^{*}$ for mechanically equivalent particles is
$\zeta^{*}=\frac{\sqrt{2}\pi^{\frac{d-1}{2}}}{d\Gamma\left(\frac{d}{2}\right)}\chi\left(1-\alpha^{2}-2\Delta^{*2}-\sqrt{2\pi}\alpha\Delta^{*}\right).$ (110)
In the steady state, the cooling rate vanishes. Thus, according to Eq. (110), the condition $\zeta^{*}=0$ yields a quadratic equation in $\Delta^{*}$ whose physical solution (i.e., $\Delta^{*}=0$ if $\alpha=1$) provides the $\alpha$-dependence of $\Delta^{*}$. It is given by
$\Delta^{*}(\alpha)=\frac{1}{2}\sqrt{\frac{\pi}{2}}\alpha\left[\sqrt{1+\frac{4(1-\alpha^{2})}{\pi\alpha^{2}}}-1\right].$ (111)
For identical particles, we have that $\Delta^{*}=\Delta/\sqrt{2T/m}$. Thus, at given values of $\alpha$ and $\Delta$, Eq. (46) gives the value of the stationary temperature. It should be noted that the relationship (111) has been tested against MD simulations, showing excellent agreement with deviations smaller than 2 %, except for small values of the coefficient of restitution and/or high densities.Brito, Risso, and Soto (2013)